### v1

@@ 2.1 Single-species without chirality
$\frac{\partial\rho}{\partial t}=\nabla^{2}\rho-\boldsymbol{\nabla}\cdot\Biggr{(}\rho\,p(\rho)\frac{\mu}{\xi^{2}}\int_{\mathbb{R}^{2}}\boldsymbol{\hat{s}}\,\Omega\left(\frac{s}{\xi}\right)\rho(\boldsymbol{x}+\boldsymbol{s},t)\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}\Biggr{)},$ (1)
with the Laplacian term representing diffusion, and the second term on the right-hand side representing advection driven by nonlocal interactions within the population. The population at $\boldsymbol{x}$ interacts with those at each point $\boldsymbol{x}+\boldsymbol{s}$. Here, $\boldsymbol{\hat{s}}$ is the unit vector in the $\boldsymbol{s}$ direction. The magnitude of interaction depends on their separation distance, $s\equiv|\boldsymbol{s}|$, through the function $\Omega(\cdot)$. This function, called the ‘interaction kernel’, is typically non-increasing and non-negative, meaning that interactions get weaker with distance and do not change direction with distance. The kernel is normalised such that the interaction strength is parameterised solely by $\mu\in\mathbb{R}$, with $\mu>0$ encoding attraction and $\mu<0$ encoding repulsion. The characteristic interaction range is parameterised by $\xi\in\mathbb{R}_{>0}$. We include a saturation function, $p(\rho)$, which is non-negative and decreases with density, representing impeded movement due to volume-filling.
@@
$\frac{\partial\rho}{\partial t}=\nabla^{2}\rho-\boldsymbol{\nabla}\cdot\Biggr{(}\rho\,p(\rho)\frac{\mu}{\xi}\boldsymbol{\nabla}\int_{\mathbb{R}^{2}}\,\Omega\left(\frac{s}{\xi}\right)\rho(\boldsymbol{x}+\boldsymbol{s},t)\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}\Biggr{)}.$ (2)
We refer to Equation (1) as a ‘direct sensing’ model, and Equation (2) as a ‘gradient sensing’ model. In direct sensing models, advection is induced directly in response to the presence of each individual, and hence is directed parallel to the separation of the individuals, i.e., in the direction of the unit vector $\boldsymbol{\hat{s}}$. The net advection is then the vector sum (integral) of these interactions. In contrast, for gradient sensing models, advection is induced in response to the gradient of a nonlocal measure of the population density.
Broadly speaking, a direct sensing framework is more often used when interactions involve one-to-one contact, such as for membrane adhesion between cells [40, 22]. The gradient sensing framework is more commonly used to study animal swarming behaviours [32, 12]. However, both models overlap in application and qualitative behaviour, with both types being used to model cellular and ecological systems [2, 41, 46, 45, 8]. In fact, with a redefinition of $\Omega$, there is a direct equivalence between these frameworks by use of the divergence theorem, given suitable conditions at any domain boundary. Nevertheless, in line with Painter et al. [43], which has a more detailed discussion of these models, we will consider them separately, as they have quite different phenomenological interpretations. We will show that our results apply in both frameworks.
@@ 2.2 Two-species with chirality
$\boldsymbol{\underline{\underline{R}}}(\alpha)\boldsymbol{\hat{s}}\equiv\begin{pmatrix}\text{cos}\,\alpha&-\text{sin}\,\alpha\\ \text{sin}\,\alpha&\text{cos}\,\alpha\end{pmatrix}\boldsymbol{\hat{s}}=\text{cos}(\alpha)\,\boldsymbol{\hat{s}}+\text{sin}(\alpha)\,(\boldsymbol{\hat{e}_{z}}\times\boldsymbol{\hat{s}}).$ (3)
@@
$\boldsymbol{\nabla}\int_{\mathbb{R}^{2}}\,\Omega\left(\frac{s}{\xi}\right)\rho(\boldsymbol{x}+\boldsymbol{s},t)\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}\longrightarrow\boldsymbol{\underline{\underline{R}}}(\alpha)\left(\boldsymbol{\nabla}\int_{\mathbb{R}^{2}}\,\Omega\left(\frac{s}{\xi}\right)\rho(\boldsymbol{x}+\boldsymbol{s},t)\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}\right).$ (4)
In both cases, $-90^{\circ}<\alpha\leq 90^{\circ}$, so that the distinction between attraction and repulsion is still governed by the sign of the interaction strength, $\mu$.
@@
$\begin{split}\frac{\partial c}{\partial t}=\nabla^{2}c-\boldsymbol{\nabla}\cdot\Biggr{(}cp_{c}(c,\rho)\Biggr{[}\frac{\mu_{cc}}{\xi_{cc}^{2}}\int_{\mathbb{R}^{2}}\boldsymbol{\underline{\underline{R}}}(\alpha_{cc})\boldsymbol{\hat{s}}\,\Omega_{cc}\left(\frac{s}{\xi_{cc}}\right)g_{cc}(c(\boldsymbol{x}+\boldsymbol{s},t))\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}&\\ +\frac{\mu_{c\rho}}{\xi_{c\rho}^{2}}\int_{\mathbb{R}^{2}}\boldsymbol{\underline{\underline{R}}}(\alpha_{c\rho})\boldsymbol{\hat{s}}\,\Omega_{c\rho}\left(\frac{s}{\xi_{c\rho}}\right)g_{c\rho}(\rho(\boldsymbol{x}+\boldsymbol{s},t))\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}\Biggr{]}\Biggr{)}\\ \frac{\partial\rho}{\partial t}=D\nabla^{2}\rho-\boldsymbol{\nabla}\cdot\Biggr{(}\rho p_{\rho}(c,\rho)\Biggr{[}\frac{\mu_{\rho\rho}}{\xi_{\rho\rho}^{2}}\int_{\mathbb{R}^{2}}\boldsymbol{\underline{\underline{R}}}(\alpha_{\rho\rho})\boldsymbol{\hat{s}}\,\Omega_{\rho\rho}\left(\frac{s}{\xi_{\rho\rho}}\right)g_{\rho\rho}(\rho(\boldsymbol{x}+\boldsymbol{s},t))\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}&\\ +\frac{\mu_{\rho c}}{\xi_{\rho c}^{2}}\int_{\mathbb{R}^{2}}\boldsymbol{\underline{\underline{R}}}(\alpha_{\rho c})\boldsymbol{\hat{s}}\,\Omega_{\rho c}\left(\frac{s}{\xi_{\rho c}}\right)g_{\rho c}(c(\boldsymbol{x}+\boldsymbol{s},t))\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}\Biggr{]}\Biggr{)},\end{split}$ (5)
where $D$ is the ratio of diffusivity between the two species. For further generality, instead of assuming that the magnitude each advection is simply directly proportional to the densities $c(\boldsymbol{x}+\boldsymbol{s},t)$ or $\rho(\boldsymbol{x}+\boldsymbol{s},t)$, we allow them to vary as some non-negative, non-decreasing functions $g_{uv}(\cdot)$.
@@
$\begin{split}\frac{\partial c}{\partial t}=\nabla^{2}c-\boldsymbol{\nabla}\cdot\Biggr{(}cp_{c}(c,\rho)\Biggr{[}\frac{\mu_{cc}}{\xi_{cc}}\,\boldsymbol{\underline{\underline{R}}}(\alpha_{cc})\boldsymbol{\nabla}\int_{\mathbb{R}^{2}}\Omega_{cc}\left(\frac{s}{\xi_{cc}}\right)g_{cc}(c(\boldsymbol{x}+\boldsymbol{s},t))\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}&\\ +\frac{\mu_{c\rho}}{\xi_{c\rho}}\,\boldsymbol{\underline{\underline{R}}}(\alpha_{c\rho})\boldsymbol{\nabla}\int_{\mathbb{R}^{2}}\Omega_{c\rho}\left(\frac{s}{\xi_{c\rho}}\right)g_{c\rho}(\rho(\boldsymbol{x}+\boldsymbol{s},t))\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}\Biggr{]}\Biggr{)}\\ \frac{\partial\rho}{\partial t}=D\nabla^{2}\rho-\boldsymbol{\nabla}\cdot\Biggr{(}\rho p_{\rho}(c,\rho)\Biggr{[}\frac{\mu_{\rho\rho}}{\xi_{\rho\rho}}\,\boldsymbol{\underline{\underline{R}}}(\alpha_{\rho\rho})\boldsymbol{\nabla}\int_{\mathbb{R}^{2}}\Omega_{\rho\rho}\left(\frac{s}{\xi_{\rho\rho}}\right)g_{\rho\rho}(\rho(\boldsymbol{x}+\boldsymbol{s},t))\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}&\\ +\frac{\mu_{\rho c}}{\xi_{\rho c}}\,\boldsymbol{\underline{\underline{R}}}(\alpha_{\rho c})\boldsymbol{\nabla}\int_{\mathbb{R}^{2}}\Omega_{\rho c}\left(\frac{s}{\xi_{\rho c}}\right)g_{\rho c}(c(\boldsymbol{x}+\boldsymbol{s},t))\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}\Biggr{]}\Biggr{)}.\end{split}$ (6)
@@ 3.1 Only the parallel component of advection affects linear stability
To find the conditions for pattern formation in these models, we perform a linear stability analysis, starting with the single species direct sensing model defined by Equation (1) and including the chiral rotation matrix defined by Equation (3). Our approach follows the method presented in Section 3 of Jewell et al. [22]. Firstly, we linearise about the homogeneous steady-state, $P$, which is determined from the total mass of the system, dictated by the initial conditions. We take $\rho(\boldsymbol{x},t)=P+\tilde{\rho}(\boldsymbol{x},t)$. This gives
$\frac{\partial\tilde{\rho}}{\partial t}=\nabla^{2}\tilde{\rho}-Pp(P)\frac{\partial g}{\partial\rho}\Big{\rvert}_{P}\,\frac{\mu}{\xi^{2}}\boldsymbol{\nabla}\cdot\left(\boldsymbol{\underline{\underline{R}}}(\alpha)\int_{\mathbb{R}^{2}}\boldsymbol{\hat{s}}\,\Omega\left(\frac{s}{\xi}\right)\tilde{\rho}(\boldsymbol{x}+\boldsymbol{s},t)\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}\right).$ (7)
@@
$\begin{split}\lambda(\boldsymbol{k})&=-k^{2}-Pp(P)\frac{\partial g}{\partial\rho}\Big{\rvert}_{P}\,\frac{\mu}{\xi^{2}}\,ik\boldsymbol{\hat{k}}\cdot\boldsymbol{\underline{\underline{R}}}(\alpha)\int_{\mathbb{R}^{2}}\boldsymbol{\hat{s}}\,\Omega\left(\frac{s}{\xi}\right)\,e^{i\boldsymbol{k}\cdot\boldsymbol{s}}\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}\\ &=-k^{2}+2\pi\,Pp(P)\frac{\partial g}{\partial\rho}\Big{\rvert}_{P}\,\frac{\mu}{\xi^{2}}\,k\,(\boldsymbol{\hat{k}}\cdot\boldsymbol{\underline{\underline{R}}}(\alpha)\boldsymbol{\hat{k}})\int_{0}^{\infty}\Omega\left(\frac{s}{\xi}\right)\,s\,J_{1}(ks)\mathop{}\!\mathrm{d}s,\end{split}$ (8)
@@
Crucially, Eq. (8) differs from the dispersion relation of the non-chiral model only through the inclusion of the factor $\boldsymbol{\hat{k}}\cdot\boldsymbol{\underline{\underline{R}}}(\alpha)\boldsymbol{\hat{k}}$. This simple result holds for any constant matrix $\boldsymbol{\underline{\underline{R}}}$, making it useful for the linear analysis of models with various forms of spatial anisotropy. For instance, in certain biological systems, movement may be more favourable in the $y$ direction than in the $x$ direction, which could be encoded by a non-identity diagonal matrix. In the case of chirality, $\boldsymbol{\underline{\underline{R}}}$ corresponds to a rotation matrix, and thus
@@
$\frac{\partial\tilde{\rho}}{\partial t}=\nabla^{2}\tilde{\rho}-Pp(P)\frac{\partial g}{\partial\rho}\Big{\rvert}_{P}\,\frac{\mu}{\xi}\boldsymbol{\nabla}\cdot\left(\boldsymbol{\underline{\underline{R}}}(\alpha)\boldsymbol{\nabla}\int_{\mathbb{R}^{2}}\Omega\left(\frac{s}{\xi}\right)\tilde{\rho}(\boldsymbol{x}+\boldsymbol{s},t)\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}\right).$ (10)
@@
$\lambda(\boldsymbol{k})=-k^{2}+2\pi\,Pp(P)\frac{\partial g}{\partial\rho}\Big{\rvert}_{P}\,\frac{\mu}{\xi}\,k^{2}\,(\boldsymbol{\hat{k}}\cdot\boldsymbol{\underline{\underline{R}}}(\alpha)\boldsymbol{\hat{k}})\int_{0}^{\infty}\Omega\left(\frac{s}{\xi}\right)\,s\,J_{0}(ks)\mathop{}\!\mathrm{d}s,$ (11)

### v2

@@ 2.1 Single-species without chirality
$\frac{\partial\rho}{\partial t}=\nabla^{2}\rho-\boldsymbol{\nabla}\cdot\Biggr(\rho\,\phi(\rho)\frac{\mu}{\xi^{2}}\int_{\mathbb{R}^{2}}\boldsymbol{\hat{s}}\,\Omega\left(\frac{s}{\xi}\right)g(\rho(\boldsymbol{x}+\boldsymbol{s},t))\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}\Biggr),$ (1)
with the Laplacian term representing diffusion, and the second term on the right-hand side representing advection driven by nonlocal interactions within the population. The population at $\boldsymbol{x}$ interacts with those at each point $\boldsymbol{x}+\boldsymbol{s}$. Here, $\boldsymbol{\hat{s}}$ is the unit vector in the $\boldsymbol{s}$ direction. For generality, instead of assuming that the advection is simply directly proportional to the density, $\rho(\boldsymbol{x}+\boldsymbol{s},t)$, we allow it to vary as some non-negative, non-decreasing function $g(\rho)$. The magnitude of interaction also depends on the separation distance, $s\equiv|\boldsymbol{s}|$, through the function $\Omega(\cdot)$. This function, called the ‘interaction kernel’, is typically non-increasing and non-negative, meaning that interactions get weaker with distance and do not change direction with distance. The kernel is normalised such that the interaction strength is parameterised solely by $\mu\in\mathbb{R}$, with $\mu>0$ encoding attraction and $\mu<0$ encoding repulsion. The characteristic interaction range is parameterised by $\xi\in\mathbb{R}_{>0}$. We include a saturation function, $\phi(\rho)$, which is non-negative and decreases with density, representing impeded movement due to volume-filling.
@@
$\frac{\partial\rho}{\partial t}=\nabla^{2}\rho-\boldsymbol{\nabla}\cdot\Biggr(\rho\,\phi(\rho)\frac{\mu}{\xi}\boldsymbol{\nabla}\int_{\mathbb{R}^{2}}\,\tilde{\Omega}\left(\frac{s}{\xi}\right)g(\rho(\boldsymbol{x}+\boldsymbol{s},t))\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}\Biggr),$ (2)
where $\tilde{\Omega}$ is the interaction kernel for this model. We refer to Equation (1) as a ‘direct sensing’ model, and Equation (2) as a ‘gradient sensing’ model. In direct sensing models, advection is induced directly in response to the presence of each individual, and hence is directed parallel to the separation of the individuals, i.e., in the direction of the unit vector $\boldsymbol{\hat{s}}$. The net advection is then the vector sum (integral) of these interactions. In contrast, for gradient sensing models, advection is induced in response to the gradient of a nonlocal measure of the population density.
Broadly speaking, a direct sensing framework is more often used when interactions involve one-to-one contact, such as for membrane adhesion between cells [40, 22]. The gradient sensing framework is more commonly used to study animal swarming behaviours [32, 12]. However, both models overlap in application and qualitative behaviour, with both types being used to model cellular and ecological systems [2, 41, 46, 45, 8]. In fact, if $\Omega=\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}s}\tilde{\Omega}$, then there is a direct equivalence between these frameworks, which can be seen by use of the divergence theorem, given suitable conditions at any domain boundary [43]. Nevertheless, in line with Painter et al. [43], which has a more detailed discussion of these models, we will consider them separately, as they have quite different phenomenological interpretations. We will show that our results apply in both frameworks.
@@ 2.2 Two-species with chirality
$\boldsymbol{\underline{\underline{R}}}(\alpha)\boldsymbol{\hat{s}}\equiv\begin{pmatrix}\text{cos}\,\alpha&-\text{sin}\,\alpha\\ \text{sin}\,\alpha&\text{cos}\,\alpha\end{pmatrix}\boldsymbol{\hat{s}}\,$. (3)
@@
$\boldsymbol{\nabla}\int_{\mathbb{R}^{2}}\,\tilde{\Omega}\left(\frac{s}{\xi}\right)g(\rho(\boldsymbol{x}+\boldsymbol{s},t))\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}\longrightarrow\boldsymbol{\underline{\underline{R}}}(\alpha)\left(\boldsymbol{\nabla}\int_{\mathbb{R}^{2}}\,\tilde{\Omega}\left(\frac{s}{\xi}\right)g(\rho(\boldsymbol{x}+\boldsymbol{s},t))\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}\right).$ (4)
In both cases, $90^{\circ}<\alpha\leq 90^{\circ}$, so that the distinction between attraction and repulsion is still governed by the sign of the interaction strength, $\mu$.
@@
$\begin{split}\frac{\partial c}{\partial t}=\nabla^{2}c-\boldsymbol{\nabla}\cdot\Biggr(c\phi_{c}(c,\rho)\Biggr[\frac{\mu_{cc}}{\xi_{cc}^{2}}\int_{\mathbb{R}^{2}}\boldsymbol{\underline{\underline{R}}}(\alpha_{cc})\boldsymbol{\hat{s}}\,\Omega_{cc}\left(\frac{s}{\xi_{cc}}\right)g_{cc}(c(\boldsymbol{x}+\boldsymbol{s},t))\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}&\\ +\frac{\mu_{c\rho}}{\xi_{c\rho}^{2}}\int_{\mathbb{R}^{2}}\boldsymbol{\underline{\underline{R}}}(\alpha_{c\rho})\boldsymbol{\hat{s}}\,\Omega_{c\rho}\left(\frac{s}{\xi_{c\rho}}\right)g_{c\rho}(\rho(\boldsymbol{x}+\boldsymbol{s},t))\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}\Biggr]\Biggr)\\ \frac{\partial\rho}{\partial t}=D\nabla^{2}\rho-\boldsymbol{\nabla}\cdot\Biggr(\rho\phi_{\rho}(c,\rho)\Biggr[\frac{\mu_{\rho\rho}}{\xi_{\rho\rho}^{2}}\int_{\mathbb{R}^{2}}\boldsymbol{\underline{\underline{R}}}(\alpha_{\rho\rho})\boldsymbol{\hat{s}}\,\Omega_{\rho\rho}\left(\frac{s}{\xi_{\rho\rho}}\right)g_{\rho\rho}(\rho(\boldsymbol{x}+\boldsymbol{s},t))\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}&\\ +\frac{\mu_{\rho c}}{\xi_{\rho c}^{2}}\int_{\mathbb{R}^{2}}\boldsymbol{\underline{\underline{R}}}(\alpha_{\rho c})\boldsymbol{\hat{s}}\,\Omega_{\rho c}\left(\frac{s}{\xi_{\rho c}}\right)g_{\rho c}(c(\boldsymbol{x}+\boldsymbol{s},t))\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}\Biggr]\Biggr),\end{split}$ (5)
where $D$ is the ratio of diffusivity between the two species.
@@
$\begin{split}\frac{\partial c}{\partial t}=\nabla^{2}c-\boldsymbol{\nabla}\cdot\Biggr(c\phi_{c}(c,\rho)\Biggr[\frac{\mu_{cc}}{\xi_{cc}}\,\boldsymbol{\underline{\underline{R}}}(\alpha_{cc})\boldsymbol{\nabla}\int_{\mathbb{R}^{2}}\tilde{\Omega}_{cc}\left(\frac{s}{\xi_{cc}}\right)g_{cc}(c(\boldsymbol{x}+\boldsymbol{s},t))\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}&\\ +\frac{\mu_{c\rho}}{\xi_{c\rho}}\,\boldsymbol{\underline{\underline{R}}}(\alpha_{c\rho})\boldsymbol{\nabla}\int_{\mathbb{R}^{2}}\tilde{\Omega}_{c\rho}\left(\frac{s}{\xi_{c\rho}}\right)g_{c\rho}(\rho(\boldsymbol{x}+\boldsymbol{s},t))\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}\Biggr]\Biggr)\\ \frac{\partial\rho}{\partial t}=D\nabla^{2}\rho-\boldsymbol{\nabla}\cdot\Biggr(\rho\phi_{\rho}(c,\rho)\Biggr[\frac{\mu_{\rho\rho}}{\xi_{\rho\rho}}\,\boldsymbol{\underline{\underline{R}}}(\alpha_{\rho\rho})\boldsymbol{\nabla}\int_{\mathbb{R}^{2}}\tilde{\Omega}_{\rho\rho}\left(\frac{s}{\xi_{\rho\rho}}\right)g_{\rho\rho}(\rho(\boldsymbol{x}+\boldsymbol{s},t))\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}&\\ +\frac{\mu_{\rho c}}{\xi_{\rho c}}\,\boldsymbol{\underline{\underline{R}}}(\alpha_{\rho c})\boldsymbol{\nabla}\int_{\mathbb{R}^{2}}\tilde{\Omega}_{\rho c}\left(\frac{s}{\xi_{\rho c}}\right)g_{\rho c}(c(\boldsymbol{x}+\boldsymbol{s},t))\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}\Biggr]\Biggr).\end{split}$ (6)
@@ 3.1 Only the parallel component of advection affects linear stability
To find the conditions for pattern formation in these models, we perform a linear stability analysis, starting with the single species direct sensing model defined by Equation (1) and including the chiral rotation matrix defined by Equation (3). Our approach follows the method presented in Section 3 of Jewell et al. [22]. Firstly, we linearise about the homogeneous steady-state, $P$, which is determined from the total mass of the system, dictated by the initial conditions. We take $\rho(\boldsymbol{x},t)=P+\tilde{\rho}(\boldsymbol{x},t)$, where $|\tilde{\rho}(\boldsymbol{x},t)|\ll 1$ is a small heterogeneous perturbation. Retaining only leading order terms yields
$\frac{\partial\tilde{\rho}}{\partial t}=\nabla^{2}\tilde{\rho}-P\phi(P)\frac{\partial g}{\partial\rho}\Big\rvert_{P}\,\frac{\mu}{\xi^{2}}\boldsymbol{\nabla}\cdot\left(\boldsymbol{\underline{\underline{R}}}(\alpha)\int_{\mathbb{R}^{2}}\boldsymbol{\hat{s}}\,\Omega\left(\frac{s}{\xi}\right)\tilde{\rho}(\boldsymbol{x}+\boldsymbol{s},t)\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}\right).$ (7)
@@
$\begin{split}\lambda(\boldsymbol{k})&=-k^{2}-P\phi(P)\frac{\partial g}{\partial\rho}\Big\rvert_{P}\,\frac{\mu}{\xi^{2}}\,ik\boldsymbol{\hat{k}}\cdot\boldsymbol{\underline{\underline{R}}}(\alpha)\int_{\mathbb{R}^{2}}\boldsymbol{\hat{s}}\,\Omega\left(\frac{s}{\xi}\right)\,e^{i\boldsymbol{k}\cdot\boldsymbol{s}}\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}\\ &=-k^{2}+2\pi\,P\phi(P)\frac{\partial g}{\partial\rho}\Big\rvert_{P}\,\frac{\mu}{\xi^{2}}\,k\,(\boldsymbol{\hat{k}}\cdot\boldsymbol{\underline{\underline{R}}}(\alpha)\boldsymbol{\hat{k}})\int_{0}^{\infty}\Omega\left(\frac{s}{\xi}\right)\,s\,J_{1}(ks)\mathop{}\!\mathrm{d}s,\end{split}$ (8)
@@
Crucially, Equation (8) differs from the dispersion relation of the non-chiral model only through the inclusion of the factor $\boldsymbol{\hat{k}}\cdot\boldsymbol{\underline{\underline{R}}}(\alpha)\boldsymbol{\hat{k}}$. This simple result holds for any constant matrix $\boldsymbol{\underline{\underline{R}}}$, making it useful for the linear analysis of models with various forms of spatial anisotropy. For instance, in certain biological systems, movement may be more favourable in the $y$ direction than in the $x$ direction, which could be encoded by a non-identity diagonal matrix. In the case of chirality, $\boldsymbol{\underline{\underline{R}}}$ corresponds to a rotation matrix, and thus
@@
$\frac{\partial\tilde{\rho}}{\partial t}=\nabla^{2}\tilde{\rho}-P\phi(P)\frac{\partial g}{\partial\rho}\Big\rvert_{P}\,\frac{\mu}{\xi}\boldsymbol{\nabla}\cdot\left(\boldsymbol{\underline{\underline{R}}}(\alpha)\boldsymbol{\nabla}\int_{\mathbb{R}^{2}}\tilde{\Omega}\left(\frac{s}{\xi}\right)\tilde{\rho}(\boldsymbol{x}+\boldsymbol{s},t)\mathop{}\!\mathrm{d}^{2}\boldsymbol{s}\right).$ (10)
@@
$\lambda(\boldsymbol{k})=-k^{2}+2\pi\,P\phi(P)\frac{\partial g}{\partial\rho}\Big\rvert_{P}\,\frac{\mu}{\xi}\,k^{2}\,(\boldsymbol{\hat{k}}\cdot\boldsymbol{\underline{\underline{R}}}(\alpha)\boldsymbol{\hat{k}})\int_{0}^{\infty}\tilde{\Omega}\left(\frac{s}{\xi}\right)\,s\,J_{0}(ks)\mathop{}\!\mathrm{d}s,$ (11)